OpenStudy (anonymous):

Simplify the following expression. a + 3c − 4b − c + a − 2b Select one: A. 2a + 6b + 2c B. a − 6b + 2c C. 2a − 6b + 2c D. 2a − 5b + 2c

OpenStudy (tcp2315):

C

OpenStudy (tcp2315):

a+a = 2a -4b-2b= -6b 3c-c=2c

OpenStudy (anonymous):

dangi thought d lol i was wrong